# Dynamics of feedback Ising model

### v1

@@ Abstract
We study the dynamics of a mean-field Ising model whose coupling depends on the magnetization via a linear feedback function. A key feature of this linear feedback Ising model (FIM) is the possibility of temperature-induced bistability, where a temperature increase can favor bistability between two phases. We show that the linear FIM provides a minimal model for a transcritical bifurcation as the temperature varies. Moreover, there can be two or three critical temperatures when the external magnetic field is non-negative. In the bistable region, we identify a Maxwell temperature where the two phases are equally probable, and we find that increasing the temperature favors the lower phase. We show that the probability distribution becomes non-Gaussian on certain time intervals when the magnetization converges algebraically at either zero temperature or critical temperatures. Near critical points in the parameter space, we derive a Fokker-Planck equation, construct the families of equilibrium distributions, and formulate scaling laws for transition rates between two stable equilibria. The linear FIM provides considerable freedom to control steady-state bifurcations and their associated equilibrium distributions, which can be desirable for modeling feedback systems across disciplines.
@@ I Introduction
The Ising model (IM), proposed as a toy model of ferromagnetism Lenz (1920); Ising (1925), acquired a paradigmatic status Baxter (1982); Itzykson and Drouffe (1989) and remains the go-to framework for explaining how simple, local interactions give rise to collective order and phase transitions in complex systems. In the classical version of the IM, the interaction strengths are fixed; the magnetic field is also imposed externally, albeit it can vary with time. Therefore, the spins do not influence the parameters that govern their behavior. Real-world complex systems are rarely so one-way. Ecological populations can modify their habitats, e.g., forests influence soil composition and local climate. Regional geography can also modify global climate, e.g., sea ice cover alters the reflectivity of the planet. In social systems, agents continually reshape the information environment they inhabit. These kinds of feedback loops, in which a macroscopic observable feeds back into the microscopic rules, can dramatically change the stability of states and transitions between states.
@@
Prior attempts to construct a feedback loop in the IM typically incorporate feedback into the external field or the temperature. Bornholdt’s market model reproduces market observations by letting the local external field depend on both the local spin and the global magnetization Bornholdt (2001), while a later variant lets the “social temperature” depend on the global magnetization Krause and Bornholdt (2012). In a different context, linear feedback endows mean-field Landau theory with an Andronov-Hopf bifurcation, which implies limit cycles in a fully connected Ising system Martino (2019). In Ref. Ma et al. (2025), we have introduced the general FIM whose coupling is an arbitrary function of the magnetization. At zero temperature, the FIM can describe binary phase transformations with an arbitrary bifurcation diagram, thereby establishing the relevance of feedback coupling to data-driven modeling. The simplest FIM with linear feedback coupling already exhibits many salient features of the general FIM and should find most applications.
@@
We limit ourselves with linear dependence of the coupling on $m$, yielding the linear FIM with the Hamiltonian
@@
Here $h$ is an external field. The feedback parameter $\gamma$ measures the strength of the feedback. Albeit positive $\gamma$ reinforces the majority orientation, while negative $\gamma$ weakens it, these two scenarios are equivalent by flipping the signs of $s_{i}$, $h$, and $\gamma$ in the Hamiltonian. This observation allows us to pose $\gamma>0$.
Our goal is to reveal the non-equilibrium properties of the FIM via Glauber dynamics, although we note that the equilibrium properties of related systems have been studied in various settings. The feedback coupling in the general FIM, e.g., $1+\gamma m]$ in the linear FIM, can be written as a Taylor series in $m$. Thus, the Hamiltonian is a weighted sum of $p$-spin interactions, $p\in\mathbb{N}$, which are called equivalent-neighbor interactions or more generally separable interactions in early literature den Ouden et al. (1976a, b). For ferromagnetic quadratic interactions, Bogolyubov Jr. Jr. (1972) used a minimax principle to derive the free energy. Den Ouden, Capel, Perk, et al. then extended this method to higher-degree separable interactions den Ouden et al. (1976a) and more generally separable functions of short-range interactions den Ouden et al. (1976b). This exact result can be simplified using a convex-envelope construction, i.e., a Legendre transform Perk et al. (1977). In this framework, the free energy is first recast from the fixed-field (canonical) ensemble, e.g., fixed-$h$, to the fixed-density ensemble, e.g., fixed-$m$, and then subject to a generalized Maxwell construction. This framework can reveal the critical behavior of systems with short-range interactions under small perturbations including critical-exponent renormalization and first-order transitions Capel et al. (1979).
@@
In contrast, early interest in 3-spin interactions largely relates to the well-known exact solution of the Baxter-Wu model Baxter and Wu (1973), i.e., the IM with 3-spin interactions on the triangular lattice. This model with additional 2-spin interactions cannot be solved exactly, but a mean-field approximation has been formulated Frøyen et al. (1976). Recently, the mean-field IM with 2-spin and 3-spin interactions has been applied to human-AI ecosystems P. et al. (2022), where the 3-spin interactions describe 2 humans interacting with an AI agent or 2 AI agents interacting with a human. The equilibrium properties of this model are further analyzed Contucci et al. (2024); Osabutey (2024). Although some of our results are known from these studies, our FIM unifies mixed $p$-spin models via feedback couplings and thus provides a general toolbox to model feedback systems. Moreover, our focus on non-equilibrium properties via Glauber dynamics aligns with popular modeling techniques from nonlinear dynamics.
This paper is organized as follows. Section II provides a warm-up by revisiting the Curie-Weiss model with Glauber single spin-flip dynamics, examining both small fluctuations around an equilibrium and thermally activated switching between two equilibria. Section III develops the FIM in full, revealing its unique features including temperature-induced bistability on the phase diagram and the dynamics of probability distributions at zero temperature. In Sec. IV, the master equation for Glauber dynamics is reduced to a Fokker-Planck (FP) equation near critical points in the parameter space, yielding asymptotic formulas for equilibrium distributions and scaling laws for transition rates. In Sec. V, we briefly discuss how such feedback-coupled interactions illuminate phenomena across magnetic, ecological, climatic, and socio-economic systems.
@@ II Curie-Weiss Model and Glauber dynamics
A classical system mimicking a phase transition goes back to Curie Curie (1895) and Weiss Weiss (1907), who tried to understand the phenomenon of spontaneous magnetization, and used what is now known as mean-field approximations. The precise setting where such approximations are exact is the IM on a complete graph, where every spin interacts with every other spin. This model is now called the Curie-Weiss model Kac (1968). This simplest exactly solvable model occupies a key place in science Thompson (2015); Friedli and Velenik (2017), and even its equilibrium characteristics continue to attract attention Kochmański et al. (2013); Barhoumi-Andréani and Eichelsbacher (2025). Endowing the Curie-Weiss model with Glauber spin-flip dynamics gives the simplest exactly solvable dynamical model that sheds light on the dynamics of the IM in finite dimensions.
@@ II.1 Zero external magnetic field
where the sum runs over all neighbors of vertex $i$. The spin-flip rate (4) is the simplest one that agrees with the detailed balance; other rates with such property exist but are analytically intractable already in one dimension Krapivsky et al. (2010). On the hypercubic lattices, the ferromagnetic IM endowed with Glauber dynamics has been solved only in one dimension Glauber (1963); Bedeaux et al. (1970); Felderhof (1971), although a lot is also known about qualitative behaviors of higher-dimensional Ising-Glauber models Krapivsky et al. (2010); Bray (1994).
@@
Below the critical temperature ($\beta>\beta_{c}$), the magnetization approaches to $\pm m_{\infty}$. From Eq. (8) we find that the approach to final magnetization is exponential:
@@
The above analysis applies to infinite systems. Let us now keep $N$ finite. The zero-temperature dynamics drives the system to a ground state with fully aligned spins. If the initial magnetization $m_{0}$ is positive, it is driven to the up phase, i.e., the number $M$ of up spins increases. Let $P_{M}(t)$ be the probability of having $M$ up spins and NMN-M down spins. At zero temperature, the probabilities $P_{M}(t)$ evolve according to
@@
The Gaussian shape can be justified by utilizing the $1/N$ expansion, while Eqs. (15a15b) can be derived from the master equation (13). The time evolution of the mean $\langle M\rangle$ in Eq. (15a) agrees with the solution of Eq. (8) with initial condition $m(t=0)=m_{0}$ at $\beta=\infty$. The variance $\Delta^{2}$ in Eq. (15b) increases from 0 at $t=0$, peaks at $t=\log 2$, and decreases to 0 as $t\to\infty$.
@@
The phase transition point in the infinite system is replaced by a finite region when the system size $N$ is finite Stanley (1971). The width of this region, known as a ‘scaling window’, vanishes as $N\to\infty$. For the Curie-Weiss model, the width of the scaling window scales as $N^{-1/2}$. To appreciate the behavior in the scaling window, one sets $\beta=1-b/\sqrt{N}$ with fixed $b$. The equilibrium distribution (17a) corresponds to $b=0$. One can generalize it Barhoumi-Andréani and Eichelsbacher (2025) to arbitrary finite $b$ and obtain the distribution
@@
$C=\sqrt{\frac{3|b|}{2}}\,e^{B}\times\begin{cases}K_{\frac{1}{4}}(B)&b>0\\ \frac{\pi}{\sqrt{2}}\left[I_{-\frac{1}{4}}(B)+I_{\frac{1}{4}}(B)\right]&b<0\end{cases}$ (19)
@@
The equilibrium distribution (17a) at the critical point is non-Gaussian, while the distribution at zero temperature, Eq. (14), is Gaussian throughout the evolution. We explore the evolution at the critical point using the FP equation for $P_{M}(t)$ derived in Appendix A. In the $t\to\infty$ limit, the solution of the FP equation is (17a); this provides an alternative derivation of the equilibrium distribution (17a). We solved the FP equation, viz., we expressed the Laplace transform of $P_{M}(t)$ via bi-confluent Heun functions Olver et al. (2010). If the initial magnetization is zero, $P_{M}(t)$ is Gaussian for $t\ll t_{c}$ and non-Gaussian for $t\gg t_{c}$; the crossover time is $t_{c}\sim\sqrt{N}$.
@@ II.2 Nonzero external magnetic field
$w_{i}=\frac{1-s_{i}\tanh(\beta(h+m))}{2}\,.$ (20)
@@
$\frac{dm}{dt}=-m+\tanh(\beta(h+m))\equiv f(\beta,m)$ (21)
@@
Figure 1: The equilibrium magnetization $m_{+},m_{0},m_{-}$, top to bottom, versus $\beta=T^{-1}$ when $h=\frac{1}{5}$. The positive equilibrium $m_{+}(\beta)>0$ exists in the entire temperature range. Two negative equilibria are created in a saddle-node bifurcation at $\beta_{c}(h)$ implicitly determined by Eq. (24). At the bifurcation point, $m_{c}(h)=m_{0}(\beta_{c})=m_{-}(\beta_{c})=-\sqrt{1-\beta_{c}^{-1}}$. In the present case, $\beta_{c}(\frac{1}{5})\approx 1.72286$ and $m_{c}(\frac{1}{5})\approx 0.647742$. In the zero-temperature limit, $m_{\pm}(\infty)=\pm 1$ and $m_{0}(\infty)=-\frac{1}{5}$; generally, $m_{0}(\infty)=-h$ when $0<h<1$.
@@
The equilibrium solution $m_{-}(\beta)$ is also defined for $\beta\geq\beta_{c}$ and decreases from $m_{c}$ to 1-1. The limiting behaviors of $m_{-}$ are
@@ II.3 Equilibrium distribution and transition time
$\begin{split}w_{-}(u)&\equiv\epsilon W_{-}(I)=\frac{1}{4}(1+u)(1-\tanh(-\beta{\cal H}^{\prime}(u)),\\ w_{+}(u)&\equiv\epsilon W_{+}(I)=\frac{1}{4}(1-u)(1+\tanh(-\beta{\cal H}^{\prime}(u)).\end{split}$ (33)
Hereinafter we often shortly write $\epsilon=\frac{1}{N}$. Replacing summation by integration we recast (32) into
$\begin{split}&-\epsilon\log(\rho_{\text{eq}}(m))+\epsilon\log(\rho_{\text{eq}}(-1))\\ =&\int_{-1}^{m}\frac{1}{2}\log\left(\frac{w_{-}(u)}{w_{+}(u)}\right)du+O(\epsilon)\\ \equiv&U(m)-U(-1)+O(\epsilon)\end{split}$ (34)

### v2

@@ Abstract
We study the dynamics of a mean-field Ising model whose coupling depends on the magnetization via a linear feedback function. A key feature of this linear feedback Ising model (FIM) is the possibility of temperature-induced bistability, where a temperature increase can favor bistability between two phases. We show that the linear FIM provides a minimal model for a transcritical bifurcation as the temperature varies. Moreover, there can be two or three critical temperatures when the external magnetic field is non-negative. In the bistable region, we identify a Maxwell temperature where the two phases are equally probable, and we find that increasing the temperature favors the lower phase. We show that the probability distribution becomes non-Gaussian on certain time intervals when the magnetization converges algebraically at either zero temperature or critical temperatures. Near critical points in the parameter space, we derive a Fokker-Planck equation, construct the families of equilibrium distributions, and formulate scaling laws for transition rates between two stable equilibria. The linear FIM offers considerable flexibility in controlling steady-state bifurcations and their associated equilibrium distributions, which can be desirable for modeling feedback systems across various disciplines.
@@ I Introduction
The Ising model (IM), proposed as a toy model of ferromagnetism Lenz (1920); Ising (1925), acquired a paradigmatic status Baxter (1982); Itzykson and Drouffe (1989) and remains the go-to framework for explaining how simple, local interactions give rise to collective order and phase transitions in complex systems. In the classical version of the IM, the interaction strengths are fixed; the magnetic field is also imposed externally, albeit it can vary with time. Therefore, the spins do not influence the parameters that govern their behavior. Real-world complex systems are rarely so one-way. Ecological populations can modify their habitats. For instance, forests influence soil composition and local climate. Regional geography can also affect global climate, e.g., sea ice cover alters the reflectivity of the planet. In social systems, agents continually reshape the information environment they inhabit. These kinds of feedback loops, in which a macroscopic observable feeds back into the microscopic rules, can dramatically change the stability of states and transitions between states.
@@
Prior attempts to construct a feedback loop in the IM typically incorporate feedback into the external field or the temperature. Bornholdt’s market model reproduces market observations by letting the local external field depend on both the local spin and the global magnetization Bornholdt (2001), while a later variant lets the “social temperature” depend on the global magnetization Krause and Bornholdt (2012). In a different context, linear feedback endows mean-field Landau theory with an Andronov-Hopf bifurcation, which implies the emergence of limit cycles in a fully connected Ising system Martino (2019). In Ref. Ma et al. (2025), we have introduced the general FIM whose coupling is an arbitrary function of the magnetization. At zero temperature, the FIM can describe binary phase transformations with an arbitrary bifurcation diagram, thereby establishing the relevance of feedback coupling to data-driven modeling. The simplest FIM with linear feedback coupling already exhibits many salient features of the general FIM and should find most applications.
@@
We limit ourselves to linear dependence of the coupling on $m$, yielding the linear FIM with the Hamiltonian
@@
Here $h$ is an external field. The feedback parameter $\gamma$ measures the strength of the feedback. Although positive $\gamma$ reinforces the majority orientation, while negative $\gamma$ weakens it, these two scenarios are equivalent by flipping the signs of $s_{i}$, $h$, and $\gamma$ in the Hamiltonian. This observation allows us to pose $\gamma>0$.
Our goal is to reveal the non-equilibrium properties of the FIM via Glauber dynamics, although we note that the equilibrium properties of related systems have been studied in various settings. The feedback coupling in the general FIM, e.g., $1+\gamma m]$ in the linear FIM, can be written as a Taylor series in $m$. Thus, the Hamiltonian is a weighted sum of $p$-spin interactions, $p\in\mathbb{N}$, which are called equivalent-neighbor interactions or, more generally, separable interactions in early literature den Ouden et al. (1976a, b). For ferromagnetic quadratic interactions, Bogolyubov Jr. Jr. (1972) used a minimax principle to derive the free energy. Den Ouden, Capel, Perk, et al. then extended this method to higher-degree separable interactions den Ouden et al. (1976a) and more generally separable functions of short-range interactions den Ouden et al. (1976b). This exact result can be simplified using a convex-envelope construction, i.e., a Legendre transform Perk et al. (1977). In this framework, the free energy is first recast from the fixed-field (canonical) ensemble, e.g., fixed-$h$, to the fixed-density ensemble, e.g., fixed-$m$, and then subject to a generalized Maxwell construction. This framework can reveal the critical behavior of systems with short-range interactions under small perturbations, including critical-exponent renormalization and first-order transitions Capel et al. (1979).
@@
In contrast, early interest in 3-spin interactions largely relates to the well-known exact solution of the Baxter-Wu model Baxter and Wu (1973), i.e., the IM with 3-spin interactions on the triangular lattice. This model with additional 2-spin interactions cannot be solved exactly; its mean-field approximation has been formulated in Ref. Frøyen et al. (1976). Recently, the mean-field IM with 2-spin and 3-spin interactions has been applied to human-AI ecosystems Contucci et al. (2022), where the 3-spin interactions describe two humans interacting with an AI agent or two AI agents interacting with a human. The equilibrium properties of this model are further analyzed Contucci et al. (2024); Osabutey (2024). Although some of our results are known from these studies, our FIM unifies mixed $p$-spin models via feedback couplings and thus provides a general toolbox to model feedback systems. Moreover, our focus on non-equilibrium properties via Glauber dynamics aligns with popular modeling techniques from nonlinear dynamics.
This paper is organized as follows. In Sec. II, we revisit the Curie-Weiss model with Glauber single spin-flip dynamics. We examine both small fluctuations around an equilibrium and thermally activated switching between two equilibria. Section III develops the FIM in full, revealing its unique features, including temperature-induced bistability on the phase diagram and the dynamics of probability distributions at zero temperature. In Sec. IV, the master equation for Glauber dynamics is reduced to a Fokker-Planck (FP) equation near critical points in the parameter space, yielding asymptotic formulas for equilibrium distributions and scaling laws for transition rates. In Sec. V, we briefly discuss how such feedback-coupled interactions illuminate phenomena across magnetic, ecological, climatic, and socio-economic systems.
@@ II Curie-Weiss Model and Glauber dynamics
A classical system mimicking a phase transition goes back to Curie Curie (1895) and Weiss Weiss (1907), who tried to understand the phenomenon of spontaneous magnetization and introduced what is now known as the mean-field approximation. The precise setting where such approximations are exact is the IM on a complete graph, where every spin interacts with every other spin. This model is now called the Curie-Weiss model Kac (1968). This simplest exactly solvable model occupies a key place in science Thompson (2015); Friedli and Velenik (2017), and even its equilibrium characteristics continue to attract attention Kochmański et al. (2013); Barhoumi-Andréani and Eichelsbacher (2025). Endowing the Curie-Weiss model with Glauber spin-flip dynamics gives the simplest exactly solvable dynamical model, shedding light on the dynamics of the IM in finite dimensions.
@@ II.1 Zero external magnetic field
where the sum runs over all neighbors of vertex $i$. The spin-flip rate (4) is the simplest one that agrees with the detailed balance; other rates with such a property exist but are analytically intractable already in one dimension Krapivsky et al. (2010). On the hypercubic lattices, the ferromagnetic IM endowed with Glauber dynamics has been solved only in one dimension Glauber (1963); Bedeaux et al. (1970); Felderhof (1971), although a lot is also known about qualitative behaviors of higher-dimensional Ising-Glauber models Krapivsky et al. (2010); Bray (1994).
@@
Below the critical temperature ($\beta>\beta_{c}$), the magnetization approaches $\pm m_{\infty}$. From Eq. (8) we find that the approach to final magnetization is exponential:
@@
The above analysis applies to infinite systems. Let us now turn to systems with large but finite $N$. The zero-temperature dynamics drive the system to a ground state with fully aligned spins. If the initial magnetization $m_{0}$ is positive, it is driven to the up phase, i.e., the number $M$ of up spins increases. Let $P_{M}(t)$ be the probability of having $M$ up spins and NMN-M down spins. At zero temperature, the probabilities $P_{M}(t)$ evolve according to
@@
The Gaussian shape can be justified by utilizing the $1/N$ expansion. Furthermore, Eqs. (15a15b) can be derived from the master equation (13). The time evolution of the mean $\langle M\rangle$ in Eq. (15a) agrees with the solution of Eq. (8) with initial condition $m(t=0)=m_{0}$ at $\beta=\infty$. The variance $\Delta^{2}$ in Eq. (15b) increases from 0 at $t=0$, peaks at $t=\log 2$, and decreases to 0 as $t\to\infty$.
@@
The phase transition point in the infinite system is replaced by a finite region when the system size $N$ is finite Stanley (1971). The width of this region, known as a ‘scaling window’, vanishes as $N\to\infty$. For the Curie-Weiss model, the width of the scaling window scales as $N^{-1/2}$. To appreciate the behavior in the scaling window, one sets $\beta=1-b/\sqrt{N}$ with $b=O(1)$. The equilibrium distribution (17a) corresponds to $b=0$. One can generalize it Barhoumi-Andréani and Eichelsbacher (2025) to arbitrary finite $b$ and obtain the distribution
@@
$C(b)=\frac{\pi\sqrt{3|b|}}{2}\,e^{B}\times\begin{cases}\frac{\sqrt{2}}{\pi}\,K_{\frac{1}{4}}(B)&b>0\\ I_{-\frac{1}{4}}(B)+I_{\frac{1}{4}}(B)&b<0\end{cases}$ (19)
@@
The equilibrium distribution (17a) at the critical point is non-Gaussian, while the distribution at zero temperature, Eq. (14), is Gaussian throughout the evolution. We explore the evolution at the critical point using the FP equation for $P_{M}(t)$ derived in Appendix A. In the $t\to\infty$ limit, the solution of the FP equation is (17a); this provides an alternative derivation of the equilibrium distribution (17a). We solved the FP equation. More precisely, we expressed the Laplace transform of $P_{M}(t)$ via bi-confluent Heun functions Olver et al. (2010). If the initial magnetization is zero, $P_{M}(t)$ is Gaussian for $t\ll t_{c}$ and non-Gaussian for $t\gg t_{c}$; the crossover time is $t_{c}\sim\sqrt{N}$.
@@ II.2 Nonzero external magnetic field
$w_{i}=\frac{1-s_{i}\tanh[\beta(h+m)]}{2}\,.$ (20)
@@
$\frac{dm}{dt}=-m+\tanh[\beta(h+m)]\equiv f(\beta,m)$ (21)
@@
Figure 1: The equilibrium magnetization $m_{+},m_{0},m_{-}$, top to bottom, versus $\beta=T^{-1}$ when $h=\frac{1}{5}$. The positive equilibrium $m_{+}(\beta)>0$ exists in the entire temperature range. Two negative equilibria are created in a saddle-node bifurcation at $\beta_{c}(h)$ implicitly determined by Eq. (24). At the bifurcation point, $m_{c}(h)=m_{0}(\beta_{c})=m_{-}(\beta_{c})=-\sqrt{1-\beta_{c}^{-1}}$. In the present case, $\beta_{c}(\frac{1}{5})\approx 1.72286$ and $m_{c}(\frac{1}{5})\approx-0.647742$. In the zero-temperature limit, $m_{\pm}(\infty)=\pm 1$ and $m_{0}(\infty)=-\frac{1}{5}$; generally, $m_{0}(\infty)=-h$ when $0<h<1$.
@@
The equilibrium solution $m_{-}(\beta)$ is also defined in the range $\beta\geq\beta_{c}$. One observes that $m_{-}(\beta)$ decreases from $m_{c}$ to 1-1 as $\beta$ increases from $\beta_{c}$ to infinity. The limiting behaviors of $m_{-}$ are
@@ II.3 Equilibrium distribution and transition time
$\begin{split}w_{-}(u)&\equiv\epsilon W_{-}(I)=\frac{1}{4}(1+u)(1-\tanh[-\beta{\cal H}^{\prime}(u)]),\\ w_{+}(u)&\equiv\epsilon W_{+}(I)=\frac{1}{4}(1-u)(1+\tanh[-\beta{\cal H}^{\prime}(u)]),\end{split}$ (33)
with $\epsilon=\frac{1}{N}$. Replacing summation by integration we recast (32) into
$\begin{split}&-\epsilon\log(\rho_{\text{eq}}(m))+\epsilon\log(\rho_{\text{eq}}(-1))\\ &=\int_{-1}^{m}\frac{1}{2}\log\left(\frac{w_{-}(u)}{w_{+}(u)}\right)du+O(\epsilon)\\ &\equiv U(m)-U(-1)+O(\epsilon)\end{split}$ (34)